# StreamingTalker: Audio-driven 3D Facial Animation with Autoregressive Diffusion Model

### v2

@@ Acknowledgments
Ethics Statements
As our approach enables realistic synthesis of facial motion from audio, we acknowledge potential risks of misuse, such as in generating deceptive or unauthorized content. We emphasize that our work is intended for legitimate applications, including virtual communication, accessibility tools, and creative production.
@@ A. Dataset Details
BIWI Dataset.  In our experiments, the dataset is divided into a training set (BIWI-Train) comprising 192 sentences spoken by six participants following (fan2022faceformer), with each participant contributing 32 sentences. The validation set (BIWI-Val) includes 24 sentences from the same six participants, with each speaking four sentences. Additionally, there are two testing sets: BIWI-Test-A, which contains 24 sentences from six participants seen during training, and BIWI-Test-B, which includes 32 sentences from eight participants not seen during training. BIWI-Test-A is suitable for quantitative and qualitative evaluations, while BIWI-Test-B is primarily used for qualitative assessment.
VOCA Dataset.  For consistency and fair comparison, we adopt the same training (VOCA-Train), validation (VOCA-Val), and testing (VOCA-Test) splits as used in VOCA (cudeiro2019voca) and (fan2022faceformer).
@@ B. Implementation Details
Network architecture.  In our experiments on the VOCASET dataset, we use a six-layer transformer architecture with eight attention heads for both the encoder and decoder of the VQ-VAE. The feature dimension is set to 1024 and the feedforward network dimension is 1536, while the size of the codebook is 256, with the number of face components set to 16. For the AR Condition Predictor, we use a two-layer transformer decoder with four attention heads, whose hidden size is 768 and the feedforward network size is 1024 while the history length is set to 60. Furthermore, The diffusion head is implemented as a single-layer MLP with the same hidden size as the AR Condition Predictor. We adopt a scaled linear schedule (ho2020denoising) for the diffusion process, with a total of $N=1000$ steps. In the denoising phase, we employ the Denoising Diffusion Implicit Models (DDIM) (song2020ddim), using only 50 steps to achieve efficient denoising while maintaining high-quality results.
For experiments on the more complex BIWI dataset, which contains more intricate data with greater upper-face movements, we scale up the model. Specifically, we increase the size of the codebook to 512, and the number of face components is set to 8 to mitigate the risk of overfitting (xing2023codetalker). The AR Condition Predictor is also scaled accordingly, with a hidden size of 1024 and a feedforward network size of 2048 while the history length is set to 120. In addition, we fine-tune the VAE decoder to better capture the rich facial expressions present in the BIWI dataset. All other configurations remain consistent with those used for the VOCASET dataset.
Training.  Our model was developed using PyTorch and trained on a single NVIDIA RTX 4090 GPU. We set the batch size to 1 and train 400 epochs using the AdamW (loshchilov2017adamw) optimizer with a learning rate initialized to 1e-4 in both stages. In the first stage of training, we decrease the learning rate by half every 20 epochs in VOCASET and 40 epochs in BIWI. The overall training takes about 12 hours.
@@ C. Real-time Demo
The client is responsible for capturing the user’s audio input and converting it to text. To achieve this, we use a local Vosk (vosk_api) model, which performs speech-to-text (STT) processing. Once the audio is transcribed into text, the client sends the resulting text to the server via a Socket (socket) connection for further processing.
@@
The server side of the demo is composed of two key components: our autoregressive AR diffusion model and a large language model (LLM). Upon receiving the text from the client, the server first utilizes the LLM to generate a relevant response based on the input text. Next, the server converts this response into speech using the Google Text-to-Speech (gtts) tool. This text-to-speech conversion allows the system to produce an audio response aligned with the generated text.
To optimize the real-time performance of the system, we employ multithreading during the model inference stage. In the main thread, the model processes the audio using the Wav2Vec2 (baevski2020wav2vec) processor, which handles the pre-processing of the speech input. Following this, the AR diffusion model performs streaming generation, producing a sequence of frames that represent the facial animation corresponding to the audio. Each generated frame is added to a global queue, allowing the frames to be processed in parallel. The render thread continuously retrieves frames from the queue and visualizes them in real time. This ensures the facial animation remains synchronized with the input speech and responds dynamically to the model’s output. To ensure the visual and audio components are properly synchronized, we implement a dedicated speech thread. This thread starts playing the audio once the first frame has been rendered and outputted by the render thread. By managing the audio playback this way, we maintain perfect synchronization between the animation and the corresponding speech, enhancing the user experience during the demo.

### v3

@@ A. Dataset Details
BIWI Dataset.  In our experiments, the dataset is divided into a training set (BIWI-Train) comprising 192 sentences spoken by six participants following (Fan et al. 2022), with each participant contributing 32 sentences. The validation set (BIWI-Val) includes 24 sentences from the same six participants, with each speaking four sentences. Additionally, there are two testing sets: BIWI-Test-A, which contains 24 sentences from six participants seen during training, and BIWI-Test-B, which includes 32 sentences from eight participants not seen during training. BIWI-Test-A is suitable for quantitative and qualitative evaluations, while BIWI-Test-B is primarily used for qualitative assessment.
VOCA Dataset.  For consistency and fair comparison, we adopt the same training (VOCA-Train), validation (VOCA-Val), and testing (VOCA-Test) splits as used in VOCA (Cudeiro et al. 2019) and (Fan et al. 2022).
@@ B. Implementation Details
Network architecture.  In our experiments on the VOCASET dataset, we use a six-layer transformer architecture with eight attention heads for both the encoder and decoder of the VQ-VAE. The feature dimension is set to 1024 and the feedforward network dimension is 1536, while the size of the codebook is 256, with the number of face components set to 16. For the AR Condition Predictor, we use a two-layer transformer decoder with four attention heads, whose hidden size is 768 and the feedforward network size is 1024 while the history length is set to 60. Furthermore, The diffusion head is implemented as a single-layer MLP with the same hidden size as the AR Condition Predictor. We adopt a scaled linear schedule (Ho et al. 2020) for the diffusion process, with a total of $N=1000$ steps. In the denoising phase, we employ the Denoising Diffusion Implicit Models (DDIM) (Song et al. 2020), using only 50 steps to achieve efficient denoising while maintaining high-quality results.
For experiments on the more complex BIWI dataset, which contains more intricate data with greater upper-face movements, we scale up the model. Specifically, we increase the size of the codebook to 512, and the number of face components is set to 8 to mitigate the risk of overfitting (Xing et al. 2023). The AR Condition Predictor is also scaled accordingly, with a hidden size of 1024 and a feedforward network size of 2048 while the history length is set to 120. In addition, we fine-tune the VAE decoder to better capture the rich facial expressions present in the BIWI dataset. All other configurations remain consistent with those used for the VOCASET dataset.
Training.  Our model was developed using PyTorch and trained on a single NVIDIA RTX 4090 GPU. We set the batch size to 1 and train 400 epochs using the AdamW (Loshchilov et al. 2017) optimizer with a learning rate initialized to 1e-4 in both stages. In the first stage of training, we decrease the learning rate by half every 20 epochs in VOCASET and 40 epochs in BIWI. The overall training takes about 12 hours.
@@ C. Real-time Demo
The client is responsible for capturing the user’s audio input and converting it to text. To achieve this, we use a local Vosk (Shmyrev 2023) model, which performs speech-to-text (STT) processing. Once the audio is transcribed into text, the client sends the resulting text to the server via a Socket (Kurose and Ross 2017) connection for further processing.
@@
The server side of the demo is composed of two key components: our autoregressive AR diffusion model and a large language model (LLM). Upon receiving the text from the client, the server first utilizes the LLM to generate a relevant response based on the input text. Next, the server converts this response into speech using the Google Text-to-Speech (pndurette 2021) tool. This text-to-speech conversion allows the system to produce an audio response aligned with the generated text.
To optimize the real-time performance of the system, we employ multithreading during the model inference stage. In the main thread, the model processes the audio using the Wav2Vec2 (Baevski et al. 2020) processor, which handles the pre-processing of the speech input. Following this, the AR diffusion model performs streaming generation, producing a sequence of frames that represent the facial animation corresponding to the audio. Each generated frame is added to a global queue, allowing the frames to be processed in parallel. The render thread continuously retrieves frames from the queue and visualizes them in real time. This ensures the facial animation remains synchronized with the input speech and responds dynamically to the model’s output. To ensure the visual and audio components are properly synchronized, we implement a dedicated speech thread. This thread starts playing the audio once the first frame has been rendered and outputted by the render thread. By managing the audio playback this way, we maintain perfect synchronization between the animation and the corresponding speech, enhancing the user experience during the demo.